### v1

# Optimization of deterministic photonic graph state generation via local operations

Abstract
Realizing photonic graph states, crucial in various quantum protocols, is challenging due to the absence of deterministic entangling gates in linear optics. To address this, emitter qubits are leveraged to establish and transfer the entanglement to photons. We introduce an optimization method for such protocols based on the local Clifford equivalency of states and the graph-shape correlated generation cost parameters. Employing this method, we achieve a 50% reduction in use of the 2-qubit gates for generation of the repeater graph states and a 65% reduction in the total gate count for 15-node random dense graphs.
Introduction.—Graph states are a family of multi-qubit states where the entanglement structure between qubit nodes is characterized by the edges of a graph [1, 2]. Photonic graph states, with photons as graph nodes, are essential resources for the realization of measurement-based [3, 4] and fusion-based [5] fault-tolerant quantum computing as well as various quantum communication protocols [6, 7, 8, 9], including quantum repeaters. Photons are unique among qubit architectures as they are mobile qubits, a feature necessary for many of the mentioned applications, and are relatively inexpensive to create and manipulate, with long coherence times.
While graph states are efficient to describe in theory [10], realizing them on photonic platforms is challenging due to the lack of photon-photon interactions in linear optics [11]. The main approaches used to-date for indirect establishment of entanglement between photons, i.e., creating graph edges, are (i) probabilistic approaches utilizing a sequence of the so-called “fusion gates”, based on photon interference, measurement, and post-selection [12, 13], and (ii) deterministic approaches utilizing a set of interacting quantum emitters such as quantum dots or nitrogen-vacancy centers [14] as ancillary qubits to induce entanglement between the photons as they are emitted [15, 16]. Between the two approaches, the former faces an exponential scaling of resource overhead and requires indistinguishable photons, while the latter is experimentally challenging due to limitations on the coherent control and the entanglement of quantum emitters.
Figure 1: Schematic circuit for sequential generation of photonic graph states. The three main components are (i) single qubit gates on photons (yellow), (ii) emitter blocks representing arbitrary operations on the set of emitter qubits between emission events (green) , and (iii) emission events (blue) corresponding to a two-qubit emitter-photon gate.
Because of better scalability and fueled by technological advancements, deterministic approaches have recently gained prominence in the community, as seen in experimental breakthroughs such as 10-qubit linear cluster states reported in quantum dot systems [17] and 14-qubit GHZ states generated by a single memory atom in a cavity [18]. However, presently achievable graph sizes, still markedly limited by experimental factors [19, 17, 20], remain far from the sizes required by many practical applications. Resource-efficient protocols are thus crucial for on-demand access to a broader, more practical range of graphs.
Deterministic methods use emitter qubits to generate, maintain, and distribute entanglement between photons as they are emitted. Figure 1 shows a quantum circuit representation for such protocols. The emission process, i.e. pumping an emitter followed by the emission of a photon, can be modeled as a 2-qubit gate, e.g., a cnot [16], between a non-existing photon (modeled by a qubit in |$\ket{0}$) and an emitter, and is always the first gate applied on each photonic register. For systems designed to be deterministic, the main constraint in the model is not allowing multi-qubit gates on the emitted photons.
Devising the necessary operations between emission events for generating a target state is not trivial. As qubit connectivity is heavily constrained, this task is fundamentally different from a general quantum circuit synthesis problem of establishing a desired quantum channel [21, 22]. To date, only a limited number of resource-conscious generation protocols have been proposed, with the majority designed for specific categories of graphs [16, 23, 24, 25] and only a few for arbitrary states [26, 27], with the optimization efforts primarily concerned with minimizing the number of emitters. Although the emitter count is a significant factor, there are other important metrics that can be considered in the circuit design, such as the required emitter coherence times and the number of two-qubit gates.
In this paper, we propose a method to reduce the generation costs of photonic graph states, by leveraging the concept of local Clifford (LC) equivalency of graphs [28]. While LC equivalency has been previously employed to optimize the preparation of graph states [22, 29, 30], this is the first time that it is applied to the case of deterministic photonic-state generation, differing notably from non-photonic cases. The method works through identifying a set of alternative intermediate graphs which can be converted to the final target state with minimal overhead, providing a selection of generation routes to choose from. Moreover, we adopt a novel approach by employing graph theoretical metrics to correlate the structural characteristics of a graph with the cost parameters of its generating circuit, such as circuit size and depth. This is used to determine which intermediate graphs are more favorable for the process, filtered solely based on their shape. The use of LC equivalency, combined with the correlations between graph shape and circuit resources, forms the basis of a versatile optimization framework for photonic state generation.
We demonstrate that a correlation between the circuit size and the edge count within the local equivalency classes of dense random graphs leads to average reductions of 65% and 46% in circuit gate count and ‘emitter depth’ (to be defined below), respectively, for 15-node graphs. In another case, an analysis of the repeater graph states (RGS), crucial resources for quantum communication protocols [9], resulted in the identification of the optimized elementary circuit blocks for the generation of an arbitrarily large RGS with a 50% reduction in the number of required two-qubit gates between emitters. Moreover, a polynomial time algorithm to generate the full local equivalency class of RGS is presented for the first time.
Background and preliminaries.—To introduce an unambiguous definition for the generation cost and the resources to optimize for, we consider the quantum circuit framework in Fig. 1. The circuit models emitter qubits as entanglement mediators to produce a desired photonic state, taking a sequential block form with each block indicated by a photon emission event. Between successive emissions, arbitrary operations are allowed on the emitter space, but only single-qubit unitaries can be applied on the emitted photons. Such a configuration, with enough emitters, has been proved to be sufficient for on-demand generation of arbitrary graph states [15]. To further simplify the cost analysis, the domain of operations considered in each step can be restricted to Clifford group unitaries and measurements in the computational basis, as they are proven to be universal for the preparation of graph states while being efficient to simulate classically [10, 31].
It should be noted that although the photonic gates can be applied in a sequential manner in each step, for each photonic register, one can combine all single qubit gates throughout the circuit into a single unitary to be applied at once. Therefore, the photonic part of the circuit has minimal contribution to its size. Besides, the photonic gates are usually realized through linear optical elements, e.g., waveplates, and thus are easier to implement compared to the gates applied on the emitter qubits. Consequently, the emitter part of the circuit plays the primary role in cost considerations.
Notable circuit parameters to consider in a cost analysis are (i) the number of emitters, (ii) the number of gates of each type, and (iii) the circuit depth. In particular, using more emitters necessitates coherent control over a larger set of qubits, and a wider extent of emitter-emitter couplings, which is challenging in practice. This emitter coupling cost can be directly reflected in the number of two-qubit gates which, apart from being difficult to implement, are also a primary source of infidelity in the circuit. The depth of the circuit is associated with the maximum number of successive gates applied on the qubits and is also indicative of the time it takes to complete the protocol; two significant factors in circuit performance. Since an emitter can undergo a reset after each measurement, as seen in protocols outlined in Refs. [27, 24], and its coherence time can be considered to reset as well, a more physically relevant metric would be the maximum number of consecutive gates on an emitter in the interval between each reset to |$\ket{0}$ and a subsequent measurement. Here we define this new metric as ‘emitter depth’, complementary to the regular circuit depth.
In our analysis, apart from the emitter count, the focus is on other experimentally significant metrics that have been mostly overlooked in the existing literature on deterministic photonic state generation, e.g., the number of emitter-emitter cnots, and the emitter depth. Nevertheless, the optimization process can be easily adapted to incorporate other customized cost functions.
Optimization framework.—Our optimization method is based on two main notions: the local equivalency of states, and correlations between the structure of a graph, which is quantified by graph theoretical metrics, and the parameters of its generating circuit.
Two quantum states are locally equivalent if they can be converted into each other by local operations only, and are termed as locally Clifford (LC) equivalent if the operations are limited to single-qubit Clifford unitaries. It has been shown [28, 32] that two graph states |$\ket{G_{1}}$ and |$\ket{G_{2}}$ are LC equivalent, if and only if their respective graphs can be mapped into each other via the successive application of a graph transformation rule known as local complementation (LC rule) [33]. When the LC rule is applied on node a in a graph $G$, the subgraph comprised of the neighbourhood $N_{a}$ of a is complemented while other edges in the graph remain intact (see Fig. 2).
Figure 2: Example of graph transformation via LC rule application on the blue node. The intermediate stage depicts the edge removal (dashed) and creation (solid) processes.
Checking for the local equivalency of two graphs and finding a sequence of LC operations connecting them can be efficiently performed [34]. Convertibility under local complementation forms a valid equivalence relation on the graphs and the equivalency class for any graph, also called an orbit, can be obtained by successive application of the LC rule on its nodes [28, 32, 35].
The LC orbit can be leveraged for cost reduction whenever a deterministic graph-to-circuit map, such as the protocol in Ref. [27], is employed to find a generation circuit for a given graph. Since any two graphs within the same orbit can be converted to one another by local unitaries, i.e., single-photon gates, and given that–as previously argued–such gates have no significant contribution to our circuit cost, the minimum generation cost for all orbit members is identical and is determined by the orbit member that yields the best circuit under the utilized graph-to-circuit mapping protocol. The goal is thus to identify and generate a more favorable alternative graph in the orbit and then convert back to the original target.
Another set of alternative targets can also be obtained by permuting the temporal emission order of the photons which provides a natural way of labelling the nodes in a graph. All isomorphic graphs found by changing the photon emission order share the same entanglement structure and thus can be considered identical once generated, conditioned by photon indistinguishability. The generation cost, however, varies from case to case as generating circuits depend on emission order [27].
To compile an exhaustive array of alternative target graphs, the LC orbit is first identified via iterative application of the LC rule to graph nodes until no new graphs can be found. Subsequently, all possible emission orderings are considered for each orbit member. The set of alternative targets is then fed to a graph-to-circuit map in order to create a list of generating circuits, which are then analyzed to find the best option(s) with respect to a chosen performance metric. For our choice of graph-to-circuit map, we make use of the protocol outlined in Ref. [27]. This mapping protocol guarantees the use of a minimal number of emitters when emission order is fixed, and runs in polynomial time, making it a suitable benchmark for evaluating cost reduction performance of our proposed method. Hereafter, comparisons are made between the optimized and original cases with the ‘original’ circuit obtained using this protocol for a given graph.
Results.—In general, finding the exhaustive set of alternative graphs is not always feasible except for small-sized or highly symmetric cases. This is because the LC orbit can grow rapidly with the size of the graph; for instance, for linear and cycle graphs the scaling is exponential [33]. In fact, determining the size of the orbit of an arbitrary graph is a #P-complete problem [36]. Due to this complexity, previous investigations into the LC orbit have been limited to graphs with a maximum of 12 nodes [22, 35]. Furthermore, the number of isomorphic graphs found by permutation of the emission order scales factorially with size, making the exhaustive search impractical for large graphs.
Here, an exhaustive analysis is conducted, considering both local equivalency and isomorphism classes of graphs with up to 7 qubits [37], enumerating the 45 entanglement classes outlined in Ref. [32]. For each class the most efficient generating circuits are identified, providing recipes for the generation of any of the $\sim 1.9\times 10^{6}$ labelled connected graphs falling under this category [38]. All the analyses presented here and below were conducted using our Python package “GraphiQ” [39], developed for simulating, manipulating, and evaluating photonic graph states and their generating circuits, run on a 3.22 GHz, 8-core, Apple M1 Pro chip with 16 GB of RAM.
While a full orbit analysis is generally impractical, sampling it is feasible by applying LC operations on a random sequence of nodes. This sample can then be used to identify correlations between graphs and their generating circuits, which can serve as guidelines for an oriented search within the orbit to reach lower-cost graphs. We remark that the type and degree of the correlations are specific to each individual orbit and also depend on the selected graph-to-circuit mapping protocol.
Figure 3: Original (solid) and optimized (dashed) cost parameters for different graph sizes. Error bars are one standard deviation for values averaged over 1000 random graphs.
As a demonstration, the correlation between the graph edge count and its circuit size was investigated for an orbit sample of size 100 for random Erdös-Rényi graphs [40] with $n$=15 nodes and $0.95$ edge probability. Averaging over 1000 instances, the Pearson correlation coefficient [41]–a number between $-1$ to $1$ with $1$ indicating a positive linear correlation–was found to be $0.83\pm 0.10$ for the number of unitaries and $0.73\pm 0.14$ for the emitter depth, both serving as circuit size indicators. Such correlations can then be employed to reduce the generation cost; specifically, we utilized an efficient edge reduction strategy (see Supplemental Material [42]) via LC operations to actively aim for the graphs in the orbit that are more likely to have a lower generation cost. For graphs with $n$=15, reductions of $65\%\pm 11$% in the number of unitaries and $46\%\pm 18$% in emitter depth were obtained by averaging over 1000 cases. Fig. 3 shows the cost reduction achieved when extending the use of this strategy to larger sizes.
Figure 4: (Top) Original RGS vs the best alternative shape in its orbit. (Bottom) The elementary circuit blocks associated with the above graphs; the implementation of the block adds two photons to the state under preparation. The alternative block features a reduced number of cnots. $U$ and $U^{\prime}$ are Clifford unitaries $HPHPX$ and $HPHP$, respectively.
Compared to a random search, the correlation-based approach does not require analysis of all the graphs and their circuits in a sample. Instead, it leverages an oriented navigation of the orbit, enabling a substantial reduction in optimization runtime. Specifically, when comparing a random search method to the use of an edge reduction strategy for graph sizes in the range of $n$=30 to 60, we observed a one to two orders of magnitude reduction in runtime while maintaining a similar performance. Refer to the Supplemental Material [42] for a more thorough discussion of this procedure.
We remark that for structured graphs, i.e, graphs with shapes that can be parameterized by size, it is possible to perform a full orbit analysis for smaller graphs to gain insight and identify general patterns that may be applicable to the larger-sized versions of them. One example of interest for such cases is the RGS. An RGS performs as an all-photonic quantum repeater and is of significant importance in facilitating the realization of long distance quantum communications [24, 43, 44]. By exploring the orbit of small cases of RGS, through a dynamical system mapping, we achieved a 50% reduction in the number of emitter-emitter cnots and were able to generalize this result for the RGS with arbitrary $n$ qubits (see Fig. 4), i.e., reducing the cnot count from $n-3$ to $n/2-2$. Specifically, as shown in the correlation diagram of Fig. 5, we define the variables of the dynamical system as the edge-count of a graph (x-axis) and the cnot count of its corresponding circuit (y-axis) for all graphs in the orbit. By tracking the changes within the correlation diagram when LC transformations were applied, we were able to not only fully navigate the known orbit of a small RGS, but also to generalize the pattern and develop, for the first time, an efficient, polynomial-time algorithm to generate the orbit of an arbitrarily large RGS; the algorithm is available in the Supplemental Material [42]. It is noteworthy that the LC orbit identification task via brute-force search methods is not an efficient and scalable process in general. With a full knowledge of the orbit, we can identify the generalized graph shape associated with the least generation cost (e.g., here the number of cnots) and determine its elementary circuit block, to be employed to lower the overall generation cost of an RGS of any size. As evident in Fig. 4, the optimal block uses one cnot gate to add two photons to the state whereas in the original case, two cnots are required, enabling the change of scaling from $n$ to $n/2$. This optimized scaling of two-qubit gates, identified systematically here, aligns with the scaling of the ad-hoc protocol described in Ref. [24], which was specifically designed for the optimal generation of RGS using a different approach and a different emission sequence. Additional information on the RGS orbit can be found in the Supplemental Material [42].
Figure 5: Number of cnots vs Number of edges correlation diagram of 100-qubit RGS. Points correspond to graphs in the orbit, with the red point being the original RGS. The inset shows the portion of the orbit with low edge counts.
Conclusion.—We introduce an optimization framework tailored for the deterministic generation of photonic graph states, a task that poses significant theoretical and experimental challenges due to its inherent constraints. Our approach employs local complementation to effectively identify alternative graphs that can serve as more resource-efficient avenues for reaching a target graph.
In our correlation analyses, we establish that specific graph metrics can be linked to some cost parameters, enabling us to manage the use of resources at an orbit search level; a 50% reduction in the requisite number of emitter-emitter gates showcases one of the potential applications of this method. Yet, the door is open for exploring other graphs of interest with more elaborate graph theoretical metrics.
It is worth noting that this optimization approach is independent of the graph-to-circuit mapping protocol employed, adding a flexible layer of applicability across existing state generation protocols. Moreover, the employed cost metrics can be tailored to meet the requirements of specific experimental platforms, e.g., as we have done with the ‘emitter depth’ to address the limitations in emitters’ coherence time. Additionally, the application can be easily extended to the probabilistic fusion-based generation protocols, as in Ref. [5], by facilitating the expansion of both the size scale and diversity of resource states used as building blocks.
In summary, our optimization framework presents a versatile method to identify resource-efficient graph state preparation schemes, paving the way for future advancements in quantum communication and computing technologies.
Acknowledgement
We thank Sophia Economou for helpful discussions. This work was supported in part by MITACS, and the Air Force Office of Scientific Research (AFOSR) under Grant FA9550-22-1-0062. Hoi-Kwong Lo acknowledges research support from NSERC and Innovative Solutions Canada.
Supplemental Materials
A: Exhaustive analysis of small graphs
To showcase the results of the exhaustive analysis, we look into an entanglement class associated with the graph (A) in Fig. S1, i.e., all graph state that are LC equivalent to this graph up to relabeling of the nodes. We have identified multiple intermediate target graphs to utilize in the generation process of any member of this entanglement class, each tailored for optimizing a certain cost parameter. In Fig. S1, examples of a case with the least number of cnots (B), a case with the smallest emitter depth (C), and a case with minimal number of unitary gates (D) are presented. More details for each option are listed in Table S1. We remark that there might be more than one best performing intermediate graphs depending on the cost function used. One can find the maximum improvement for each cost parameter by comparing the best values with the worst ones present in the last column, showing the maximum of the cost parameters over this class when no optimization is done. Overall, any arbitrary cost function can be employed to filter the results of such an exhaustive analysis, e.g, the fidelity of the final state with the target when a proper noise model is employed. The full dataset containing the results from the analysis of all 45 entanglement classes of graph with $n\leq 7$ nodes is available at Ref. [1].
Figure S1: A set of optimal intermediate graphs (B-D) for generation of the entanglement class associated with graph (A), minimizing different cost functions, in the same order from (B) to (D): cnot count, emitter depth, and unitary gate count. There might be more than one best graph for each metric in the orbit. Node labels indicate the photon emission order.
Table S1: Summary of the cost metrics for the alternative graphs outlined in Fig. S1.
    Metrics Graph $A$ Graph $B$ Graph $C$ Graph $D$ Orbit Max
# Emitters $2$ $2$ $3$ $2$ $3$
# cnots 3 2 3 2 10
Depth [^1] Emitter depth as defined in the main text 14 23 7 10 41
# Unitaries 19 48 18 16 96
B: Correlation analysis
In a correlation analysis of the local equivalency class of a graph, the goal is to determine whether there exists a significant correlation between a graph metric (a parameter related to the shape of a graph) and a cost parameter of its generating circuit. We examine how changes in the average values of one metric (chosen as the dependent variable) correspond to variations in the other metric (chosen as the independent variable) across the graphs in the orbit or in a sample from it. For instance, in Fig. S2, the independent variable (x-axis) the is number of edges in the graph and the dependent variable (y-axis) is chosen as the average of the circuit gate-count over the graphs sharing the same number of edges in the sample. This diagram is derived for a sample of the LC orbit of a random graph of size $n$=20 with an edge probability of $p$=0.95. The Pearson correlation coefficient [2] for this specific diagram is $94\%$. Repeating the same procedure for 100 random graphs of the same type ($n$=20, $p$=0.95), we find the distribution of the correlation coefficients to be as seen in Fig. S3. This shows an average Pearson correlation of $0.86\pm 0.07$, indicating that for most of the graphs of this type, a strong correlation (over 0.5) holds true.
Figure S2: The correlation diagram for the average number of unitary gates in a circuit vs the number of edges, over a sample of size 200 from the orbit of the graph. The error bars indicate one standard deviation when more than one graph is associated with the same number of edge in the sample.
Figure S3: The distribution of Pearson coefficients for ‘edge count vs gate count’ correlations for 100 trials with different random graphs of size 20 and edge probability 0.95
With the correlation confirmed, one can then, in this case, aim for the graphs in the orbit with a smaller number of edges, corresponding to circuits with fewer total number of unitary gates. To do this, we developed an active edge reduction strategy to get to the sparser graphs in the orbit through applying local complementation operations on specific nodes, starting from the initial graph. In particular, we select the node with the highest product of node degree (number of neighbors) and clustering coefficient [3], and apply local complementation on it. The clustering coefficient of a node is defined as the ratio of the number of existing edges in the neighborhood of that node, to the maximum possible number of edges in that neighborhood. The process is iteratively repeated on the resulting graph until further steps would no longer reduce the number of edges.
This active edge reduction strategy shows a one to two order of magnitude optimization speed up compared with a random search method, where a random sample from the graphs in the orbit is taken, converted into corresponding circuits, and the best circuit (with minimum unitary gate count) is selected. Fig. S4 shows a runtime scaling diagram for graphs of size n=30 to 55 while the sample size is chosen so that the performance of the both methods is similar as seen in Fig. S5.
Figure S4: Runtime comparison for the same performance between the active edge reduction method (red) and a random search method (green). Error bars are 1 standard deviation for data averaged over 100 trials.
Figure S5: Unitary gate count for optimization with the active edge reduction strategy (red), the random search method (green), compared to the initial gate count(black). Error bars are 1 standard deviation for data averaged over 100 trials.
C: Repeater graph state analysis
Figure S6: Pattern in RGS orbit leading to identification of elementary circuit blocks to realize different shapes. From left to right, in each step a pair of arms from the RGS is folded into the central highly connected component. Circuit block (a) can be used to add a regular arm (a pair of photons: a core node and a leave attached to it) to the graph under preparation, while the circuit block (b) adds a folded arm (a pair of photons: two core nodes connected to all other core nodes except to each other) to the graph. The unitary gate $U$ is $HPHPX$. The number of cnot gates will be reduced as a consequence of using block (b) instead of (a) when increasing the size of the graph.
The correlation analysis results in identifying interesting patterns in the case of repeater graph states. As evident in the Fig. 5 of main text, a subclass of the orbit shows almost perfect inverse linear correlation between the cnot and edge counts. By studying this subclass of the orbit, one can find the sequence of LC operations required so that the graph, and its corresponding circuit structure, moves one step toward the lower-cost region. Such an analysis gives rise to an efficient algorithm to not only generate the correlated subclass of the orbit, but also the full LC equivalency class for any RGS with arbitrary size. The procedure is outlined in detail in Algorithm 1 below.
It is worth noting that determining the LC equivalency of a graph is not a trivial task; in fact, even determining the size of the orbit is a #P-complete problem [4]. In general as one needs to perform a depth or breadth-first search over sequences of local complementation operations on graph nodes, and check for isomorphisms of newly found graphs in each step which is itself an inefficient task. Access to the whole orbit enables one to confirm the results found through correlation analysis and moreover to perform exhaustive optimization with respect to different cost functions.
Figure S7: The best alternative graph shapes in the orbit of an $n$-node RGS with odd (left) and even (right) number of arms ($n/2$= 9 and 8 respectively) in its original form. Larger cases can be obtained by adding pairs of nodes to the graphs such that the added two nodes are connected to each other and to the two central nodes.
Figure S6 shows the pattern observed in the shape of the graphs as we move from one point to the next in the correlated part of Fig. 5. By looking into the generating circuit for each case, we were able to find different types of elementary circuit blocks that can be utilized to increase the number of photons in any part of the state. Ultimately, the best (with respect to the use of cnot gates) alternative shape in the orbit and its respective generating circuit can be identified; as depicted in Fig. S7 and Fig. S8.
Algorithm 1 RGS orbit enumeration
Define orbit as an empty list
Add the original RGS $G_{0}$ to the orbit
Define core-list as the list of indices of the core nodes in graph $G_{0}$
Apply LC transformation on one of the core nodes $i$ of $G_{0}$ to get a new graph $G$
Add $G$ to orbit
Remove $i$ from the core-list
while core-list is NOT empty do
     In $G$, apply LC transformation on one of the nodes $j$ in the core-list to get the new graph; set it as $G_{1}$
     Add $G_{1}$ to orbit
     Remove $j$ from core-list
     In $G_{1}$, apply LC transformation on the core node $i$ to get the new graph; set it as $G_{2}$
     Add $G_{2}$ to orbit
     if core-list NOT empty then
         In $G_{1}$, apply LC transformation on one of the nodes $k$ in the core-list to get the new graph; set it as $G_{3}$
         Add $G_{3}$ to orbit
         Remove $k$ from core-list
         Set $G\leftarrow G_{3}$
     end if
end while
Figure S8: The circuit to generate the best alternative graph in the orbit of an RGS with $n=2m+8$ qubits. The elementary circuit block in the middle is repeated for $1\leq k\leq m$ while the rest of the circuit is fixed. For odd $m$: $V_{1}$ = X, $V_{2}$ = U, and $V_{3}$ = $U^{\prime}$, and for even $m$: $V_{1}$ = I, $V_{2}$ = $V_{3}$ = H. $U$ and $U^{\prime}$ are Clifford unitaries $HPHPX$ and $HPHP$ respectively.